OpenStudy (anonymous):

use points (-2, 0) and (4, -1) then use slope formula to calculate slope in simplest form

OpenStudy (anonymous):

to create the line use \[\frac{ y-y1 }{ x-x1 }=\frac{ y2-y1 }{ x2-x1 }\]